Find a quadratic function that includes the set of values below.
(0,3) (2,11)(3,6)

Answer by ankor@dixie-net.com(22740)

Find a quadratic function that includes the set of values below.
(0,3) (2,11)(3,6)
Using the form ax^2 + bx + c = y
From the first set we know that c = 3
Write an equation for the last two sets
x=2; y=11
a(2^2) + b(2) + 3 = 11
4a + 2b = 11 - 3
4a + 2b = 8
simplify divide by 2
2a + b = 4
and
x=3, y=6
9a + 3b + 3 = 6
9a + 3b = 6 - 3
9a + 3b = 3
simplify divide by 3
3a + b = 1
:
Use elimination to find a
3a + b = 1
2a + b = 4
--------------subtracting eliminate b
a = -3
Find b using the equation 2a + b = 4
2(-3) + b = 4
b = 4 + 6
b = 10
:
The equation: f(x) = -3x^2 + 10x + 3
:
You can confirm this for yourself, replace x with given value, see that it = y